OpenStudy (anonymous):

In triangle ABC, angle B is 25 degrees more than 3 times angle A and angle C is 29 degrees more than twice angle A. How many degrees are in angle A?

OpenStudy (anonymous):

angle b= 3x angle c= 2x+29 3x=2x+29 x=29

OpenStudy (anonymous):

I GOT 21

OpenStudy (anonymous):

OH OOOPS

OpenStudy (anonymous):

angle b= 3x+25 angle c=2x+29 3x+25=2x+25 x+25=29 x=4

OpenStudy (anonymous):